OpenStudy (anonymous):

solve ln x= 1- ln (x+2)

OpenStudy (anonymous):

I get : lnx+ ln (x+2) ln(x)(x+2) ln (x^2 + 2x)= 1 e^1 = x^2 +2x ? and i need help from there on

zepdrix (zepdrix):

I guess we would have to complete the square on the X's. Take half of the b term, and square it.\[\large x^2+bx\]

OpenStudy (anonymous):

so x^2+2x+1 = e ?

OpenStudy (anonymous):

but i get\[x= \sqrt{e}-1\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (blurbendy):

log(x) = 1 - log(x + 2) log(x) - 1 + log(x+2) = 0 log(x(x+2)) = 1 x(x+2) = e x^2 + 2x = e x^2 + 2x + 1 = 1 + e (x + 1)^2 = 1 + e x+1 = Sqrt[1 + e] or x + 1 = - Sqrt[1 + e} x = Sqrt[1 + e] - 1 or x + 1 = - Sqrt[1 + e] x = Sqrt[1 + e] - 1 or x = -1 -Sqrt[1 + e] Substitute back into the original equation, only 1 will be right x = Sqrt[1 + e] - 1

zepdrix (zepdrix):

Ahhh sorry website froze :( erased all my stuff...

OpenStudy (anonymous):

thanks everyone and that sucks! @zepdrix :/ but thanks either way !

OpenStudy (anonymous):

ln(x)+ln(x+2)=1 ln(x(x+2))=ln(e) x(x+2)=e \[x ^{2}+2x-e=0\] D=4+4e=4(1+e) =>sqrtD=2sqrt(1+e) x1=[-2+2sqrt(1-e)]/2=-1+sqrt(1+e)>0 accepted x2=[-2-2sqrt(1-e)]/2=-1-sqrt(1+e)<0 denied